200 ml of buffer needed using stock solutions of .196M HCHOO and .271M NaCHOO. compute volume of each needed to reach PH of 4.4
Answer #1

Henderson - Hasselbalch equation is

pH = pKa + log([A-]/[HA])

4.4 = 3.75 + log( [ HCOO-]/[HCOOH])

log([HCOO-]/[HCOOH]) = 0.65

[HCOO-]/[HCOOH] = 4.467

moles of HCOO-/moles of HCOOH = 4.467

( 0.271mol/1000ml) × X / (0.196mol/ 1000ml) × ( 200ml - X) = 4.467

X 0.000271mol/ml / ( 0.0392mol - X 0.000196mol/ml) = 4.467

X0.000271mol/ml = 0.1751mol - X0.0008756mol/ml

X0.001147mol/ml = 0.1751mol

X = 152.7ml

200ml - 152.66ml = 200ml - 152.66ml = 47.3ml

Therefore,

Volume of 0.196M HCOOH required = 47.3ml

Volume of 0.271M NaHCOO required = 152.7ml